Consider an amphoteric hydroxide, M(OH)_2(s). wher

Please show all work!

Ch 13B, 22
Answer #1

pH of thesolution-7 pOH 14-pH - 14-7 now tofind OH pOH log OH - 10 1.0x10 M solubility equilibriuim equation: M(OH,(s)>M(a)+2pHof the solution 10 pOH-14-10 [OH-7-10 = 10- 1.0 10 M MOH, (s)()+20H (aq) 1.0x 10 2s IM EQM Kspof MO416 10x10 +2s 4x104 = (